OpenStudy (xfdgbh536t):

i will give medal and fan. 3,10,17,24, ... what will be the 100th number in the pattern?

alones (alones):

So this is adding by 7? Lol

OpenStudy (xfdgbh536t):

what will be the 100th number in the pattern?

OpenStudy (xfdgbh536t):

i will give medal and fan.

OpenStudy (xfdgbh536t):

help

OpenStudy (sshayer):

\[tn=a+(n-1)d\] where a=first term=3 n=100 common difference d=10-3=7 find t100=?

OpenStudy (pythagoras123):

nth term = 7n-4 7(1) - 4 = 3 7(2) - 4 = 10 7(3) - 4 = 17 . . . 7(100)-4= 696 (answer)